OpenStudy (anonymous):

a painter and his son can paint a room together in 8 hrs. if the father works aloe for 3 hours and then is joined by his son, the two together can complete the job in 6 hours. how long will it take each person alone to paint the room ?

OpenStudy (amistre64):

3f + 3(f+s) = 6 f+s = 8 is what i read from the jargon

OpenStudy (amistre64):

3f + 3(f+s) = 6 f+s = 8 3f+3f+3s = 6 f+s = 8 6f+3s = 6 f +s = 8 6f+3s = 6 -3f -3s =-24 ----------- 3f = -18 f = -6 -6 + s = 8 s = 14 doesnt make a whole lot of sense tho

OpenStudy (amistre64):

maybe 8(f+s) = 8 but im sure this is the wrong way to go about it :/

OpenStudy (amistre64):

8hrs * (f+s) = 1 job f+s = 1/8 job per hour 3hrs*f + 3*(f+s) = 1 job f + 1/8 = 1/3 f = 1/3 - 1/8 = 5jobs/24hrs; or 1job in 24/5 hrs

OpenStudy (amistre64):

sounds good, but still doesnt make sense to me quite right

OpenStudy (amistre64):

f + f + f + 1/8 + 1/8 + 1/8 = done it took 3 hours to do 3/8 of the room; which means the father did 5/8 of it in 3 hours by himself. 5/8 divided by 3 = 5/24ths in an hour ...

OpenStudy (anonymous):

@amistre64 how come my friend got father=12 son=24 ?

OpenStudy (amistre64):

dunno, youd have to ask your friend :) im not even sure if my thought was correct.

OpenStudy (amistre64):

i know the f+s is good; since working together they can do 1/8 of the job in an hour 3 hours working alone, 3 hours working with son = 6 hours makes sense to me, but it would appear that the son actually slows things down if im right